A new planet is discovered who’s mass is the same as that of the earth although it’s radius is three times smaller than earth radius find acceleration of gravity on the surface of the new planet calculate answer to one decimal
1
Expert's answer
2020-04-20T10:43:33-0400
"g'=G\frac{M}{(\frac{R}{3})^2}=9G\frac{M}{R^2}=9g"

"g'=9(9.8)=88.2\frac{m}{s^2}"